OpenStudy (anonymous):

Solve. 81^x = 27^(x + 2)

OpenStudy (lgbasallote):

81^x = 27^(x+2) (3^4)^x = (3^3)^(x+2) 3^(4x) = 3^(3x + 6) 4x = 3x + 6 x = 6